OpenStudy (anonymous):

Find a general formula for the nth term of the geometric series with: a6=1 and r=1/5

OpenStudy (jdoe0001):

\(\Large a_{\color{brown}{ n}}=a_1r^{{\color{brown}{ n}}-1}\) does that ring a bell?

OpenStudy (anonymous):

A(n) = (6^5)(1/5)^(n-1)

OpenStudy (jdoe0001):

yes no no the foggiest just posting for kicks my cat typed it in wrongly copied and pasted ?

OpenStudy (anonymous):

@jdoe0001 yea thanks

OpenStudy (anonymous):

@alyssa_nicole thats not an option :-( there is A(n) = (5^6)(1/5)^(n-1) though, is that what you meant?

OpenStudy (anonymous):

yess.

OpenStudy (jdoe0001):

ok well... we know what the 6th term is , \(\bf a_6 = 1\) and we know what "r" is, or the common ratio, \(\bf r = \frac{1}{5}\) so... what's the 1st term? well.. we dunno but we know the 6th term, and the common ratio thus